# Chen, Simchi-Levi, and Wang

### v2

@@ 5.2 Lower Bound for Experimentation with SST
6 Estimation with Perfect Information about the Control Arm
In previous sections, we demonstrated how sharing information among test arms can effectively reduce variance. In this section, we explore this concept further by considering scenarios where we have Perfect Information (PI) about the control arm. Typically, the control arm represents a default policy that may have been operational for an extended period within the system, allowing the organization to accumulate a substantial dataset and develop a thorough understanding of this baseline. Extending Example 1.1, the company might already possess ample historical data on customer behaviors, enabling it to accurately determine the customers’ transition matrices when not offering coupon. We assume the historical data accurately characterizes the customer’s behavior (the same transition) at present. In other words, there is no distribution shift between the offline dataset and the current trial. Access to extensive historical user data is a common strategy in marketing, utilized to optimize selling policies (Simester et al. 2006, Yang et al. 2024), predict future purchases (Jacobs et al. 2016), or gauge customer responses (Narayandas and Sengupta 2023). Similarly, the assumption of having an offline dataset is prevalent in other domains involving MDPs, such as offline reinforcement learning (Levine et al. 2020), reinforcement learning with human feedback (RLHF, Chang et al. 2024), and healthcare applications (Yu et al. 2021, Hu et al. 2024).
Formally, we assume that the experimenter has access to both the precise transition matrix ${\bm{P}}^{\mathfrak{c}}$ and the exact mean reward ${\bm{r}}^{\mathfrak{c}}$ of control policy $\pi^{\mathfrak{c}}$. With this perfect information, she can directly computing ${\bm{V}}^{\mathfrak{c}}$ using the Bellman equation. Consequently, her primary focus during experimental periods becomes gathering data on the new treatment policy $\pi^{\mathfrak{t}}$, applying it consistently. Moreover, since $\pi^{\mathfrak{t}}$ is SST, a substantial amount of information from ${\bm{r}}^{\mathfrak{c}}$ and ${\bm{P}}^{\mathfrak{c}}$, especially those associated with the non-crucial states, can be shared when estimating ${\bm{V}}^{\mathfrak{t}}$. More specifically, after consistently applying $\pi^{\mathfrak{t}}$, she only needs to estimate the transitions and mean reward associated with the crucial state $s^{1}$, which is
$\widehat{{\bm{P}}}^{{\mathfrak{t}}}_{\textrm{PI}}(s^{1},s)=\frac{\sum_{i=1}^{T}\mathbf{1}[s_{i}=s^{1},s_{i+1}=s]}{1\vee\sum_{i=1}^{T}\mathbf{1}[s_{i}=s^{1}]}\,\forall s\in\mathcal{S}\quad\textrm{ and }\quad\widehat{r}_{\textrm{PI}}^{\mathfrak{t}}(s^{1})=\frac{\sum_{i=1}^{T}\mathbf{1}[s_{i}=s^{1}]r_{i}}{1\vee\sum_{i=1}^{T}\mathbf{1}[s_{i}=s^{1}]}.$
For all the other components in $\widehat{{\bm{P}}}_{\textrm{PI}}^{{\mathfrak{t}}}$ and $\widehat{{\bm{r}}}_{\textrm{PI}}^{{\mathfrak{t}}}$, we directly plug in those in ${\bm{P}}^{\mathfrak{c}}$ and ${\bm{r}}^{\mathfrak{c}}$, i.e., $\widehat{{\bm{P}}}^{{\mathfrak{t}}}_{\textrm{PI}}(s,s^{\prime})={\bm{P}}^{\mathfrak{c}}(s,s^{\prime})$ and $\widehat{r}_{\textrm{PI}}^{\mathfrak{t}}(s)=r^{\mathfrak{c}}(s)$ for all $s\neq s^{1}$ and any $s^{\prime}\in\mathcal{S}$. Applying $\widehat{{\bm{P}}}_{\textrm{PI}}^{{\mathfrak{t}}}$ and $\widehat{{\bm{r}}}_{\textrm{PI}}^{{\mathfrak{t}}}$ to the Bellman equation generates an estimator $\widehat{{\bm{V}}}^{\mathfrak{t}}_{\textrm{PI}}$. Then, the difference between $\widehat{{\bm{V}}}^{\mathfrak{t}}_{\textrm{PI}}$ and ${{\bm{V}}}^{\mathfrak{c}}$ yields an estimator $\widehat{\bm{\Delta}}^{\mathrm{PI}}$ of the treatment effects. The following theorem presents the central limit theorem of our estimator. The detailed proof is delayed to Appendix 15.1.
Theorem 6.1 (Asymptotic Normality of Perfect Information)
Assuming the Markov chain induced by the SST policy $\pi^{\mathfrak{t}}$ is uniformly ergodic, it holds that as $T\to\infty$,
[…]
where
$\Sigma_{\mathrm{PI}}=({\bm{I}}-\gamma{\bm{P}}^{\mathfrak{t}})^{-1}\begin{bmatrix}\frac{\sigma_{\mathfrak{t}}^{2}(s^{1})}{\mu^{\mathfrak{t}}(s^{1})}+\frac{\left\|{\gamma{\bm{V}}^{\mathfrak{t}}}\right\|_{\Sigma_{p^{\mathfrak{t}}}(s^{1})}^{2}}{\mu^{\mathfrak{t}}(s^{1})}&\\ &{\bm{0}}\end{bmatrix}({\bm{I}}-\gamma{\bm{P}}^{\mathfrak{t}})^{-\top}\,.$
Specifically, for the crucial state $s^{1}$, it holds that as $T\to\infty$,
[…]
where
$\sigma^{2}_{\mathrm{PI}}(s^{1})=\frac{\rho^{\mathfrak{t}}_{s^{1}}(s^{1})^{2}}{\mu^{\mathfrak{t}}(s^{1})}\left({\sigma_{\mathfrak{t}}^{2}(s^{1})+\left\|{\gamma{\bm{V}}^{\mathfrak{t}}}\right\|^{2}_{\Sigma_{p^{\mathfrak{t}}}(s^{1})}}\right)\,.$
Since we are implementing a fixed policy $\pi^{\mathfrak{t}}$ during the experimental periods, it is logical that the asymptotic variance $\Sigma_{\mathrm{PI}}$ and $\sigma^{2}_{\mathrm{PI}}(s^{1})$ exhibit similar structures with $\Sigma_{\mathrm{MB}}$ and $\sigma^{2}_{\mathrm{MB}}(s^{1})$ from the model-based estimator under a fixed policy discussed in Section 3. Unlike the approach in that section, which requires estimating transitions for every state, here we only need to estimate the transitions and rewards associated with the crucial state $s^{1}$. Recalling Corollary 3.5, the asymptotic variance $\sigma^{2}_{\textrm{MB}}(s^{1})$ is the sum of variances associated with all states. With PI, $\sigma^{2}_{\textrm{PI}}(s^{1})$ only keeps the variance linked to the crucial state $s^{1}$, getting rid of all the other terms in $\sigma^{2}_{\textrm{MB}}(s^{1})$. This reduction aligns with our expectations, given the complete information on the transitions and rewards for non-crucial states now at our disposal.
We now would like to compare the variance with A/B testing, both with and without information sharing, as discussed in Sections 4 and 5. In addition to the perfect information about the control arm, as we are running $\pi^{\mathfrak{t}}$ all the time instead of the mixed policy like A/B testing, we naturally accumulate more data specific to $\pi^{\mathfrak{t}}$. Hence it’s natural to anticipate a further reduction of variance associated with crucial state $s^{1}$. To facilitate a detailed comparison and ensure consistency in our analysis, we introduce two notations, albeit redundant, $\Sigma^{\mathrm{het}}_{\mathrm{PI}}$ and $\Sigma^{\mathrm{hom}}_{\mathrm{PI}}$. These notations are designed to differentiate the parts of $\sigma^{2}_{\mathrm{PI}}(s^{1})$ related to the crucial states $s^{1}$ and all other states, respectively. Thus, $\Sigma^{\mathrm{het}}_{\mathrm{PI}}=\sigma^{2}_{\mathrm{PI}}(s^{1})$ and $\Sigma^{\mathrm{hom}}_{\mathrm{PI}}=0$ since $\sigma^{2}_{\mathrm{PI}}(s^{1})$ is only related to $s^{1}$. Formally, we have the following theorem characterizing the variance reduction of PI, whose detailed proof is delayed to Appendix 15.2.
Theorem 6.2
$\Sigma^{\mathrm{het}}_{\mathrm{PI}}<\left\{{1\land\frac{1+\delta L}{2}}\right\}\Sigma^{\mathrm{het}}_{\mathrm{AB}}\quad\textrm{and}\quad\Sigma^{\mathrm{hom}}_{\mathrm{PI}}=0\leq\Sigma^{\mathrm{hom}}_{\mathrm{IS}}\leq\Sigma^{\mathrm{hom}}_{\mathrm{AB}}\,,$
where $\delta=D_{\mathrm{TV}}({\bm{P}}^{\mathfrak{t}}(s^{1},\cdot),{\bm{P}}^{\mathfrak{c}}(s^{1},\cdot))$ and $L\coloneqq(\ln(C)+1)/(1-\lambda)$. The constants $C,\lambda$ are defined in the uniform ergodicity of the Markov chain (Definition 3.3) induced by $\pi^{1/2}$.
The parameter $\delta$ measures the difference between the transition probabilities under the treatment and the control. One of the reasons that such a parameter takes an effect is that the denominators in $\Sigma^{\mathrm{het}}_{\mathrm{AB}}$ and $\Sigma^{\mathrm{het}}_{\mathrm{PI}}$ are $\mu_{1/2}(s^{1})$ and $\mu^{{\mathfrak{t}}}(s^{1})$ respectively, which are the stationary probabilities of the crucial state $s^{1}$. A little difference in the transition probability may have an significant impact on the reciprocal of the stationary probability. Also, recall that $\Sigma^{\mathrm{het}}_{\mathrm{AB}}=\Sigma^{\mathrm{het}}_{\mathrm{IS}}$, and Theorem 6.2 confirms that PI can always strictly outperform A/B testing even with information sharing. Moreover, if the shift of transition probability $\delta$ is small or the Markov chain mixes fast enough, Theorem 6.2 indicates that PI can approximately halve the variance compared to A/B testing for the heterogeneous component. To reiterate, PI eliminates uncertainty stemming from all non-crucial states.
Similar ideas can also be generalized to SST with multiple test arms, where the perfect information about control arm is available. Following the notations in Section 5.3, assume there are $n$ test arm with the first arm serving as the control. Consequently, sampling is required only for the $n-1$ test arms excluding the control arm, such that $\pi^{1/(n-1)}(a\mid s)\coloneqq\frac{1}{n-1}\sum_{i\neq 1}\pi^{i}(a\mid s)$ for any $(s,a)$ pair. For all the test arms, we estimate only ${\bm{P}}^{{\mathfrak{t}}^{i}}(s^{1},\cdot)$ and $r^{{\mathfrak{t}}^{i}}(s^{1})$ using sample means, while inheriting the other values from the known control arm. Using the Bellman equation, we derive the estimator $\widehat{{\bm{V}}}^{{\mathfrak{t}}^{i}}$ for each SST, leading to the estimator of $\Delta_{i,j}(s^{1})$ by taking the difference. In Corollary 6.3, we establish the asymptotic normality for the estimation of $\Delta_{i,j}(s^{1})$, confirming the statistical robustness of our approach under these settings.
Corollary 6.3
Assuming the Markov chain induced by the mixed policy $\pi^{1/(n-1)},\pi^{1/n}$ are both uniformly ergodic, it holds that as $T\to\infty$,
[…]
where $\sigma_{\mathrm{PIn},(i,j)}^{2}(s^{1})$ is given in Appendix 15.1.1. Furthermore, we introduce $\Sigma^{\mathrm{het}}_{\mathrm{PIn}}(i,j)$ and $\Sigma^{\mathrm{hom}}_{\mathrm{PIn}}(i,j)$ to denote the terms in $\sigma_{\mathrm{PIn},(i,j)}^{2}(s^{1})$ associated with $s^{1}$ and all the other states respectively. We have
$\Sigma^{\mathrm{het}}_{\mathrm{PIn}}(i,j)<\left\{{1\land\frac{n-1}{n}\left({1+\frac{2\delta L}{n}}\right)}\right\}\Sigma^{\mathrm{het}}_{\mathrm{ABn}}(i,j)\quad\textrm{and}\quad\Sigma^{\mathrm{hom}}_{\mathrm{PIn}}(i,j)=0\,,$
where $\delta=\max_{i}D_{\mathrm{TV}}({\bm{P}}^{i}(s^{1},\cdot),{\bm{P}}^{1}(s^{1},\cdot))$ and $L\coloneqq(\ln(C)+1)/(1-\lambda)$. The constants $C,\lambda$ are related to the uniform ergodicity of the Markov chain induced by $\pi^{1/n}$ from Definition 3.3.
With $n$ test arms, PI continues to yield asymptotically normal estimators and significant variance reduction. Moreover, the term $\left\{{1\land\frac{n-1}{n}\left({1+\frac{2\delta L}{n}}\right)}\right\}$ is almost non-decreasing with increasing $n$. This pattern indicates that as $n$ grows, the relative benefit of variance reduction for $\Sigma^{\mathrm{het}}_{\mathrm{PIn}}(i,j)$ diminishes. Such an observation aligns with the intuition that as the number of test arms grows, the data available for each test arm decreases, potentially becoming a limiting factor in further reducing variance.
@@ 7 From SST to Local Treatment
Having demonstrated the effectiveness of IS and PI for variance reduction under SST, we can now extend these results to Local Treatment scenarios. It can be expected that all A/B testing methods—including IS and PI—under local treatment will still produce asymptotically normal estimators, similar to Theorem 4.1, 5.1 and 6.1. Compared to SST, the heterogeneous part of the system becomes a subset ${\mathcal{S}}^{\mathfrak{t}}$ instead of the single just state $s^{1}$. As a result, we can only share information, or plug in perfect information, of state $s\in{\mathcal{S}}\setminus{\mathcal{S}}^{\mathfrak{t}}$ for IS and PI estimator respectively. The following theorem characterizes this key difference from SST.
@@
Assuming the Markov Chain induced by policy $\pi^{\mathfrak{t}},\pi^{1/2}$ are both uniformly ergodic, with Local Treatment on state ${\mathcal{S}}^{\mathfrak{t}}\subseteq{\mathcal{S}}$, it holds as $T\to\infty$, for $\Gamma\in\{\mathrm{AB},\mathrm{IS},\mathrm{PI}\}$ and any state $s\in{\mathcal{S}}$,
[…]
@@
where $\Sigma_{\Gamma}^{\mathrm{het}}$ and $\Sigma_{\Gamma}^{\mathrm{hom}}$ refers to the variance of states ${\mathcal{S}}^{\mathfrak{t}}$ and ${\mathcal{S}}\setminus{\mathcal{S}}^{\mathfrak{t}}$ respectively for $\Gamma\in\{\mathrm{AB},\mathrm{IS},\mathrm{PI}\}$. The full expression of $\sigma_{\mathrm{AB}}^{2}(s),\sigma_{\mathrm{IS}}^{2}(s),\sigma_{\mathrm{PI}}^{2}(s)$ can be found in Appendix 16.
Recall in SST, $\Sigma_{\Gamma}^{\mathrm{het}}$ is associated with the single crucial state $s^{1}$, while now we generalize it to the crucial subset ${\mathcal{S}}^{\mathfrak{t}}$. For the asymptotic relative efficiency of three estimators, we then have the following generalization of Theorem 5.2 and Theorem 6.2.
@@
$\Sigma^{\mathrm{het}}_{\mathrm{IS}}=\Sigma^{\mathrm{het}}_{\mathrm{AB}}\quad\textrm{and}\quad\Sigma^{\mathrm{hom}}_{\mathrm{PI}}=0\leq\Sigma^{\mathrm{hom}}_{\mathrm{IS}}\leq\Sigma^{\mathrm{hom}}_{\mathrm{AB}}\,.$
@@
Theorem 7.2 replicates Theorem 5.2. The difference is that now we generalize $\Sigma_{\Gamma}^{\mathrm{het}}$ to variance associated with the crucial subset ${\mathcal{S}}^{\mathfrak{t}}$ rather than single crucial state $s^{1}$ in SST. This demonstrates that the effectiveness of IS diminishes as the number of states that can share information (i.e. the homogeneous part) decreases. In the extreme case of a general treatment where ${\mathcal{S}}^{\mathfrak{t}}={\mathcal{S}}$, meaning all states are treated and no states to share information, IS will reduce to A/B testing. One caveat here is that unlike Theorem 6.2, we do not have general guarantees like $\Sigma^{\mathrm{het}}_{\mathrm{PI}}<\Sigma^{\mathrm{het}}_{\mathrm{AB}}$ for local treatment. The intuition is that $\mu^{\mathfrak{t}}(s)$ can be much smaller than $\frac{1}{2}\mu_{1/2}(s)$ for some state $s$ when considering the local treatment. Hence this brings about large variance for the estimation of $\widehat{\Delta}^{\mathrm{PI}}(s)$. We leave it to future work for efficient design of experiments with perfect information and local treatment. Finally, we can derive a similar generalization to local treatment with multiple test arms as Theorem 5.6 and Corollary 6.3. We delay the details to the Appendix 16.1.

### v3

@@ 5 From SST to Local Treatment
Having demonstrated the effectiveness of IS for variance reduction under SST, we can now extend these results to Local Treatment scenarios. It can be expected that we are able to establish similar results to Theorem 3.1, 4.1. Compared to SST, the heterogeneous part of the system becomes a subset ${\mathcal{S}}^{\mathfrak{t}}$ instead of the single just state $s^{1}$. As a result, we can only share information of state $s\in{\mathcal{S}}\setminus{\mathcal{S}}^{\mathfrak{t}}$. The following theorem characterizes this extension from SST.
@@
Assuming the Markov Chain induced by policy $\pi^{\mathfrak{t}},\pi^{1/2}$ are both uniformly ergodic, with Local Treatment on state ${\mathcal{S}}^{\mathfrak{t}}\subseteq{\mathcal{S}}$, it holds as $T\to\infty$, for $\Gamma\in\{\mathrm{AB},\mathrm{IS}\}$ and any state $s\in{\mathcal{S}}$,
$\sqrt{T}\left({\widehat{\Delta}^{\Gamma}(s)-\Delta(s)}\right)\xrightarrow{d}{\mathcal{N}}(0,\sigma_{\Gamma}^{2}(s))\,$,
@@
where $\Sigma_{\Gamma}^{\mathrm{het}}$ and $\Sigma_{\Gamma}^{\mathrm{hom}}$ refers to the variance of states ${\mathcal{S}}^{\mathfrak{t}}$ and ${\mathcal{S}}\setminus{\mathcal{S}}^{\mathfrak{t}}$ respectively for $\Gamma\in\{\mathrm{AB},\mathrm{IS}\}$. The full expression of $\sigma_{\mathrm{AB}}^{2}(s),\sigma_{\mathrm{IS}}^{2}(s)$ can be found in Appendix 14.
Recall in SST, $\Sigma_{\Gamma}^{\mathrm{het}}$ is associated with the single crucial state $s^{1}$, while now we generalize it to the crucial subset ${\mathcal{S}}^{\mathfrak{t}}$. For the asymptotic relative efficiency of three estimators, we then have the following generalization of Theorem 4.2.
@@
$\Sigma^{\mathrm{het}}_{\mathrm{IS}}=\Sigma^{\mathrm{het}}_{\mathrm{AB}}\quad\textrm{and}\quad\Sigma^{\mathrm{hom}}_{\mathrm{IS}}\leq\Sigma^{\mathrm{hom}}_{\mathrm{AB}}\,.$
@@
Theorem 5.2 replicates Theorem 4.2. The difference is that now we generalize $\Sigma_{\Gamma}^{\mathrm{het}}$ to variance associated with the crucial subset ${\mathcal{S}}^{\mathfrak{t}}$ rather than single crucial state $s^{1}$ in SST. This demonstrates that the effectiveness of IS diminishes as the number of states that can share information (i.e. the homogeneous part) decreases. In the extreme case of a general treatment where ${\mathcal{S}}^{\mathfrak{t}}={\mathcal{S}}$, meaning all states are treated and no states to share information, IS will reduce to A/B testing. Finally, we can derive a similar generalization to local treatment with multiple test arms as Theorem 4.6. We delay the details to the Appendix 14.1.
6 Information Sharing Always Improves Efficiency for Differentiable Estimators
In the previous sections, we have established the optimal asymptotically unbiased estimators for local treatments in MDPs. In practice, however, decision-makers can have many other design choices. For example, a slight bias may be acceptable if it substantially reduces variance. Moreover, beyond long-term cumulative outcomes, practitioners may also be interested in other long-term treatment effects, such as average or point outcomes. This raises a fundamental question: does information sharing always improve performance regardless of the estimating approach or estimand? To address this, we introduce a broad class of estimators, differentiable estimators, which includes many widely used methods and, importantly, always benefits from information sharing.
6.1 Differentiable Estimators
We now define the differential estimators in the vanilla A/B testing without information sharing. For any piece of data we collect $X_{i}=(s_{i},a_{i},s_{i+1},r_{i})$ and $a\in\{{\mathfrak{t}},{\mathfrak{c}}\}$, we denote
${\bm{K}}^{a}_{\mathrm{AB}}(X_{i})\coloneqq 2{\bm{E}}_{s_{i},s_{i+1}}\mathbf{1}[a_{i}=a]\,,\qquad{\bm{R}}^{a}_{\mathrm{AB}}(X_{i})\coloneqq 2{\bm{e}}_{s_{i}}r_{i}\mathbf{1}[a_{i}=a]\,$,
where ${\bm{E}}_{s_{i},s_{i+1}}\in\mathbb{R}^{n\times n}$ is a matrix with all entries equal to zero except for a one at the $(s_{i},s_{i+1})$ entry and ${\bm{e}}_{s_{i}}\in\mathbb{R}^{n}$ is the standard basis vector with a one in the $s_{i}$-th position and zeros elsewhere. Using these definitions, we can have the IPW estimators for $\operatorname{Diag}({\bm{\mu}}_{1/2}){\bm{P}}^{a}$ and $\operatorname{Diag}({\bm{\mu}}_{1/2}){\bm{r}}^{a}$,
$\overline{{\bm{K}}^{a}_{\mathrm{AB}}}\coloneqq\frac{1}{T}\sum_{i=1}^{T}{\bm{K}}^{a}_{\mathrm{AB}}(X_{i})\,,\qquad\overline{{\bm{R}}^{a}_{\mathrm{AB}}}\coloneqq\frac{1}{T}\sum_{i=1}^{T}{\bm{R}}^{a}_{\mathrm{AB}}(X_{i})\,.$
Definition 6.1 (Differentiable Estimator)
An estimator $\widehat{{\bm{\Delta}}}^{\textrm{AB}}$ is called a differentiable estimator if
$\widehat{{\bm{\Delta}}}^{\textrm{AB}}=f(\overline{{\bm{K}}^{\mathfrak{t}}_{\mathrm{AB}}},\overline{{\bm{K}}^{\mathfrak{c}}_{\mathrm{AB}}},\overline{{\bm{R}}^{\mathfrak{t}}_{\mathrm{AB}}},\overline{{\bm{R}}^{\mathfrak{c}}_{\mathrm{AB}}})\,$, (6)
where $f$ is any function that is differentiable w.r.t to $\overline{{\bm{K}}^{\mathfrak{t}}_{\mathrm{AB}}},\overline{{\bm{K}}^{\mathfrak{c}}_{\mathrm{AB}}},\overline{{\bm{R}}^{\mathfrak{t}}_{\mathrm{AB}}},\overline{{\bm{R}}^{\mathfrak{c}}_{\mathrm{AB}}}$. All the differentiable estimators constitute the class $\mathcal{F}$}.
Note that although our previous ATE $\Delta$ is defined as the difference between two value functions, the estimand ${\bm{\Delta}}^{\mathrm{AB}}$ of differentiable estimator $\widehat{{\bm{\Delta}}}^{\mathrm{AB}}$ can be more general. It can also represent, for example, the difference in average rewards, the difference in finite-horizon rewards, or even the difference in advantage functions. Moreover, the function $f$ need not yield unbiased estimators, and we impose no restrictions beyond differentiability.
Furthermore, we highlight that Definition 6.1 and Eq. (6) subsumes many existing and widely used estimators for MDPs, including:
  • Naive estimator of Farias et al. (2022);
  • Difference-in-Q (DQ) estimator of Farias et al. (2022);
  • Off-policy evaluation estimator in Farias et al. (2022);
  • Longitudinal surrogate model of Huang et al. (2023);
  • Our unbiased A/B testing estimators in Algorithm 1.
This demonstrates that our framework provides a unifying perspective on several strands of the recent literature.
Remark 6.2
The fact that our model-based A/B testing estimator in Algorithm 1 belongs to $\mathcal{F}$} has deeper theoretical implications. In fact, every differentiable model-based estimator falls within $\mathcal{F}$}. More concretely, suppose the MDP is parameterized by $({\bm{P}}^{\mathfrak{t}},{\bm{P}}^{\mathfrak{c}},{\bm{r}}^{\mathfrak{t}},{\bm{r}}^{\mathfrak{c}})$. A differentiable model-based estimator first obtains the estimators $(\widehat{{\bm{P}}}^{\mathfrak{t}},\widehat{{\bm{P}}}^{\mathfrak{c}},\widehat{{\bm{r}}}^{\mathfrak{t}},\widehat{{\bm{r}}}^{\mathfrak{c}})$ via Algorithm 2, and then computes the estimate of ATE as $\widehat{{\bm{\Delta}}}^{\textrm{AB}}=g(\widehat{{\bm{P}}}^{\mathfrak{t}},\widehat{{\bm{P}}}^{\mathfrak{c}},\widehat{{\bm{r}}}^{\mathfrak{t}},\widehat{{\bm{r}}}^{\mathfrak{c}})$, where $g$ is some functional differentiable w.r.t $\widehat{{\bm{P}}}^{\mathfrak{t}},\widehat{{\bm{P}}}^{\mathfrak{c}},\widehat{{\bm{r}}}^{\mathfrak{t}},\widehat{{\bm{r}}}^{\mathfrak{c}}$. We denote the class of all differentiable model-based estimators as $\mathcal{M}$}. The following proposition establishes the inclusion $\mathcal{M}$$\subseteq{\mathcal{F}}$ by explicitly bridging $\widehat{{\bm{P}}}^{\mathfrak{t}},\widehat{{\bm{P}}}^{\mathfrak{c}},\widehat{{\bm{r}}}^{\mathfrak{t}},\widehat{{\bm{r}}}^{\mathfrak{c}}$ to $\overline{{\bm{K}}^{\mathfrak{t}}_{\mathrm{AB}}},\overline{{\bm{K}}^{\mathfrak{c}}_{\mathrm{AB}}},\overline{{\bm{R}}^{\mathfrak{t}}_{\mathrm{AB}}},\overline{{\bm{R}}^{\mathfrak{c}}_{\mathrm{AB}}}$.
Proposition 6.3
Given the trajectory $\tau=\{(s_{i},a_{i},s_{i+1},r_{i})\}_{i\in[T]}$ sampled under policy $\pi$ and assume all states are visited at least once in $\tau$, then it holds $\mathcal{M}$$\subseteq{\mathcal{F}}$.
Proof 6.4
[…]
6.2 Variance Reduction for Differentiable Estimators
Applying the information-sharing idea to differentiable estimators, we again observe a linear reduction in variance for the non-treated states, without introducing additional bias. For clarity, we focus on the SST case, and the extension to general local treatments follows naturally. In particular, information sharing modifies the estimation of model parameters at every state except $s^{1}$. In particular, for $X_{i}=(s_{i},a_{i},s_{i+1},r_{i})$ and $a\in\{{\mathfrak{t}},{\mathfrak{c}}\}$, we have
${\bm{K}}^{a}_{\mathrm{IS}}(X_{i})\coloneqq\bigg{(}2\mathbf{1}[a_{i}=a,s_{i}=s^{1}]+\mathbf{1}[s_{i}\neq s^{1}]\bigg{)}{\bm{E}}_{s_{i},s_{i+1}}\,,\quad{\bm{R}}^{a}_{\mathrm{IS}}(X_{i})\coloneqq\bigg{(}2\mathbf{1}[a_{i}=a,s_{i}=s^{1}]+\mathbf{1}[s_{i}\neq s^{1}]\bigg{)}{\bm{e}}_{s_{i}}r_{i}\,.$
And similarly, we denote $\overline{{\bm{K}}^{a}_{\mathrm{IS}}}\coloneqq\frac{1}{T}\sum_{i=1}^{T}{\bm{K}}^{a}(X_{i})$ and $\overline{{\bm{R}}^{a}_{\mathrm{IS}}}\coloneqq\frac{1}{T}\sum_{i=1}^{T}{\bm{R}}^{a}(X_{i})$. Finally, we plug in the estimators into the same $f$ and obtain our information sharing estimator,
$\widehat{{\bm{\Delta}}}^{\textrm{IS}}=f(\overline{{\bm{K}}^{\mathfrak{t}}_{\mathrm{IS}}},\overline{{\bm{K}}^{\mathfrak{c}}_{\mathrm{IS}}},\overline{{\bm{R}}^{\mathfrak{t}}_{\mathrm{IS}}},\overline{{\bm{R}}^{\mathfrak{c}}_{\mathrm{IS}}})\,.$
An immediate observation is that both $\overline{{\bm{K}}^{a}_{\mathrm{IS}}}$ and $\overline{{\bm{K}}^{a}_{\mathrm{AB}}}$ converge almost surely to $\operatorname{Diag}({\bm{\mu}}_{1/2}){\bm{P}}^{a}$ for $a\in\{{\mathfrak{t}},{\mathfrak{c}}\}$. Similarly, $\overline{{\bm{R}}^{a}_{\mathrm{IS}}}$ and $\overline{{\bm{R}}^{a}_{\mathrm{AB}}}$ converge almost surely to $\operatorname{Diag}({\bm{\mu}}_{1/2}){\bm{r}}^{a}$. Since $f$ is differentiable (and thus continuous), ${\bm{K}}^{a}_{\mathrm{AB}}(X_{i})$ and ${\bm{K}}^{a}_{\mathrm{IS}}(X_{i})$ have exactly the same asymptotic bias. More formally, and of greater importance, we have the following variance analysis.
Theorem 6.5
For any differentiable estimators $\widehat{{\bm{\Delta}}}^{\Gamma}=f(\overline{{\bm{K}}^{\mathfrak{t}}_{\Gamma}},\overline{{\bm{K}}^{\mathfrak{c}}_{\Gamma}},\overline{{\bm{R}}^{\mathfrak{t}}_{\Gamma}},\overline{{\bm{R}}^{\mathfrak{c}}_{\Gamma}})$ where $\Gamma\in\{\mathrm{AB},\mathrm{IS}\}$, $\widehat{{\bm{\Delta}}}^{\mathrm{AB}}$ and $\widehat{{\bm{\Delta}}}^{\mathrm{IS}}$ are both asymptotic normal with the same asymptotic bias. Furthermore, their asymptotic variance can be decomposed as $\Sigma_{\Gamma}=\Sigma^{\mathrm{het}}_{\Gamma}+\Sigma^{\mathrm{hom}}_{\Gamma}+\Sigma^{\mathrm{cov}}_{\Gamma}$ where
$\Sigma^{\mathrm{het}}_{\mathrm{IS}}=\Sigma^{\mathrm{het}}_{\mathrm{AB}}\,,\quad\Sigma^{\mathrm{cov}}_{\mathrm{IS}}=\Sigma^{\mathrm{cov}}_{\mathrm{AB}}\,,\textrm{ and }\quad\Sigma^{\mathrm{hom}}_{\mathrm{IS}}\preceq\Sigma^{\mathrm{hom}}_{\mathrm{AB}}.$
Following Theorem 4.2, the term $\Sigma^{\mathrm{het}}_{\Gamma}$ captures the variance associated with the non-treated states, while $\Sigma^{\mathrm{hom}}_{\Gamma}$ reflects the variance of the state targeted by the SST. The new term introduced here is $\Sigma^{\mathrm{cov}}_{\Gamma}$, which represents the covariance across time. Because the system operates under an MDP, future states depend on the past states and actions. Compared with Theorem 4.2, we can learn that selecting an appropriate functional form of $f$ can eliminate temporal correlation.
For general estimators to A/B/n testing, we can also rigorously show that information sharing can linearly reduce part of the variance. The asymptotic variance can be similarly decomposed as $\Sigma_{\Gamma}(i,j)=\Sigma^{\mathrm{het}}_{\Gamma}(i,j)+\Sigma^{\mathrm{hom}}_{\Gamma}(i,j)+\Sigma^{\mathrm{cov}}_{\Gamma}(i,j)$ where
$\Sigma^{\mathrm{het}}_{\mathrm{IS}}(i,j)=\Sigma^{\mathrm{het}}_{\mathrm{AB}}(i,j)\,,\quad\Sigma^{\mathrm{cov}}_{\mathrm{IS}}(i,j)=\Sigma^{\mathrm{cov}}_{\mathrm{AB}}(i,j)\,,\textrm{ and }\quad\Sigma^{\mathrm{hom}}_{\mathrm{IS}}(i,j)\preceq\frac{2}{n}\Sigma^{\mathrm{hom}}_{\mathrm{AB}}(i,j).$
We defer the formal statement and proofs to Appendix 12.2.